OpenStudy (josee):

Find the area of the largest rectangle that can be inscribed in a semi-circle of radius 3.

OpenStudy (anonymous):

using calculus?

OpenStudy (josee):

yepp

OpenStudy (anonymous):

ok the equation of a circle that is centered at the origin with radius 3 is x^2 + y^2 = 3^2

OpenStudy (anonymous):

if you put a rectangle in there, it will have coordinates ( -x, sqrt (9-x^2)) and (x, sqrt (9-x^2)), (x,0) and (-x,0)

OpenStudy (anonymous):

the area of the rectange is A = 2x * y , where y = sqrt (9 -x^2)

OpenStudy (anonymous):

A' = 2 sqrt (9-x^2) + 2x (1/2) ( 9-x^2)^(-1/2)(-2x)

OpenStudy (josee):

ohhh. i thought the area was 2x2y which made it 4xy

OpenStudy (anonymous):

A = 2sqrt(9-x^2) - x^2/ sqrt (9-x^2)

OpenStudy (anonymous):

common denominator that gives us A = [ 2(9-x^2) - x^2] / ... , and we dont care about denominator, so 18-3x^2 = 0

OpenStudy (anonymous):

so x = sqrt 6 is your max for x, so your base will have 2sqrt 6 , and y = sqrt 3

OpenStudy (anonymous):

so i get 12

OpenStudy (anonymous):

(l/2)^2+(w/2)^2=3^2=9 (1) and u know that A=lw for rectangle so solve equation (1) in terms of length and substitute into this area equation l=sqrt(9-(w/2^2)) so A=lw=(sqrt(9-(w/2^2))(w) then take derivative of this and set it equal to zero and solve for W which will give you maximum width dA/dw=(1/2)(9-(w/2^2))*(w^2)+(1/2)(9-(w/2^2)(w)=0 solve for w

OpenStudy (josee):

how? the answer choices are 9, 6, 5, 8, and 7 sq. units

OpenStudy (anonymous):

and once u solve for that then use equation (1) to solve for corresponding length...

OpenStudy (anonymous):

i found an error

OpenStudy (anonymous):

A = [ 2(9-x^2) -2 x^2] / sqrt 9-x^2

OpenStudy (anonymous):

so max x = 3/sqrt 2

OpenStudy (anonymous):

= 3 sqrt 2 / 2

OpenStudy (anonymous):

now we want 2x , so the base is 3 sqrt 2

OpenStudy (josee):

uhh

OpenStudy (anonymous):

josee are you boy or girl

OpenStudy (josee):

i got lost looking for w, and i dont understand your last step fermats

OpenStudy (josee):

um boy. why?

OpenStudy (anonymous):

here is the solution in a pdf http://science.kennesaw.edu/~plaval/math1190/maxmin.pdf

OpenStudy (anonymous):

scroll down

OpenStudy (anonymous):

the answer is 9

OpenStudy (anonymous):

0 Find the area of the largest rectangle that can be inscribed in a semi circle of radius r. For simplicity, we consider the upper half of the circle of radius r, centered at the origin. Its equation is x 2 + y 2 = r 2 with y 0. Thus y = p r 2 x 2 . We label (x; y) the upper right corner of the rectangle. Thus, the rectangle has height y and width 2x. Let A denote the area of this rectangle. Then, A = 2xy. We want to maximize A. First, we express A as a function of one other quantity.

OpenStudy (anonymous):

for sum reason i don't feel like the answer is 9 if it it is 9 then what did u get fermet for l and w ?

OpenStudy (anonymous):

She said x=3/sqrt(2), so w=6/sqrt(2) and h=3/sqrt(2).